Larken has forecast sales for the next three months as follows: July 4,000 units, August 6,000 units, September 7,500 units. Larken’s policy is to have an ending inventory of 40% of the next month's sales needs on hand. July 1 inventory is projected to be 1,600 units. Monthly manufacturing overhead is budgeted to be $17,000 plus $5 per unit produced. What is budgeted manufacturing overhead for July?

$41,500

$24,500

$47,000

$27,000
Answer #1

Production for July = Budgeted sales units + Desired ending finished goods units – Beginning finished goods units

= 4,000 + 6,000 x 40% - 1,600

= 4,000 + 2,400 - 1,600

= 4,800 units

Monthly manufacturing overhead is budgeted to be $17,000 plus $5 per unit produced.

Budgeted manufacturing overhead for July = 17,000 + 4,800 x 5

= 17,000 + 24,000

= $41,000

Note: There seems to be a mistake in drafting the answer choices. The correct answer is $41,000 but there is no such option given in the question. Answer choice (a) given as $41,500 must be $41,000

Kindly comment if you need further assistance. Thanks